# Weight Space Representation Learning with Neural Fields

### v1

@@ Abstract
In this work, we investigate the potential of weights to serve as effective representations, focusing on neural fields. Our key insight is that constraining the optimization space through a pre-trained base model and low-rank adaptation (LoRA) can induce structure in weight space. Across reconstruction, generation, and analysis tasks on 2D and 3D data, we find that multiplicative LoRA weights achieve high representation quality while exhibiting distinctiveness and semantic structure. When used with latent diffusion models, multiplicative LoRA weights enable higher-quality generation than existing weight-space methods.
@@ 1 Introduction
Neural network weights have traditionally been viewed as opaque byproducts of optimization, high-dimensional vectors that encode learned functions but resist interpretation or manipulation. This perspective has begun to shift with recent advances in weight space learning, where researchers have demonstrated that network parameters can be merged yang2024modelmerging; singhhyperalign; lim2024empirical, generated erkocc2023hyperdiffusion; peebles2023learning, or used as inputs to other networks zhou2023nft; hospedales2020meta; lim2024learning. A fundamental question nonetheless remains largely unexplored: Can neural network weights themselves serve as meaningful representations for data?
@@
We investigate this question in the context of implicit neural representations (INRs), where neural networks are trained to overfit individual samples by mapping coordinates to values. INRs have proven to be versatile, capable of encoding diverse data modalities within a unified architecture xie2022neural. Since INRs inherently encode signals as network parameters, using these weights as representations is a natural next step. However, neural network weights are known to be ambiguous by nature, for example because neuron permutations and scaling leave the network function unchanged zhao2025symmetry. As such, different random initializations may yield vastly different parameter configurations yet functionally identical models. In other words, functionally identical networks can be arbitrarily far in weight space zhao2025symmetry, challenging the use of network weights as data representations.
@@
Figure 1: LoRA based weight space representation with neural fields. Given an input coordinate $\mathbf{p}\in\mathbb{R}^{n}$, a base neural field is adapted via multiplicative LoRA weights $\boldsymbol{\phi}_{i}$ to produce signal values $\mathbf{v}_{\mathbf{p}}\in\mathbb{R}^{m}$, each weight representing an instance. The LoRA weights themselves form structured representations in weight space, enabling diverse applications including reconstruction, generation, and clustering.
@@
Our key insight is that constraining the network weights of different samples by introducing appropriate inductive biases allows us to transform these chaotic parameters into organized, semantic representations. To this end, as illustrated in Figure 1, we propose to use Low-Rank Adaptation (LoRA) hu2022lora within a pre-trained base neural field to create structured weight space representations. This design is motivated by two key properties of LoRA: First, adaptation through LoRA constrains the weight updates to lie within a low-dimensional subspace defined by the base model. Hu et al. hu2022lora demonstrate this through subspace similarity analysis, showing that adaptations at different ranks share common singular vector directions, suggesting the existence of a meaningful low-rank adaptation subspace. Second, the low-rank constraint inherently reduces the dimensionality of the weight space representation, mitigating the curse of dimensionality that would otherwise hinder learning in high-dimensional parameter spaces.
@@
Notably, we find that the standard additive LoRA formulation is insufficient for weight space learning in the context of neural fields. Instead, we introduce multiplicative LoRA (mLoRA), where weight updates are applied through element-wise multiplication rather than addition. This multiplicative formulation naturally aligns with modulation mechanisms in generative neural fields anokhin2021cips; karras2021alias; chan2020pigan, where features are composed through multiplicative interactions, enabling effective weight space learning.
@@
We validate our approach across multiple data modalities and tasks. First, we establish that LoRA-based weight representations achieve lower reconstruction error, greater robustness to initialization, and better linear mode connectivity than vanilla INR weights. Second, we demonstrate that these structured weight spaces support high-quality generative modeling, with diffusion models trained on multiplicative LoRA weights outperforming prior attempts in weight space neural field generation. Finally, through evaluation on discriminative tasks (i.e. classification and clustering), we confirm that the weight space structure correlates with semantic properties of the encoded data.
@@
1. We demonstrate that neural network weights, when properly constrained, can serve as effective data representations that capture semantic structure.
2. We introduce multiplicative LoRA for neural fields, which we show provides superior representation quality compared to standard additive LoRA and vanilla weight parameterizations.
3. We validate weight space representations across diverse tasks: Reconstruction, generation, and analysis, establishing their viability as a representation paradigm.
@@ 2.1 Weight Space Learning
The treatment of neural network weights as learnable representations has emerged as a distinct research direction, progressing from early hypernetwork approaches to sophisticated weight-space generative models. Early work demonstrated that network parameters could be generated by auxiliary networks ha2016hypernetworks, though these methods suffered from prohibitive memory overhead when scaling to modern architectures wang2025scaling. The fundamental challenge of weight space manipulation stems from permutation symmetry: Functionally identical networks can have vastly different parameter configurations due to neuron reordering gao2024revisiting; navon2023equivariant.
@@
Recent advances have addressed these challenges through multiple strategies. Model merging techniques leverage optimal transport and activation matching to align neurons before parameter fusion singh2020model; ormaniec2025fusion, while equivariant architectures explicitly respect weight-space symmetries when processing network parameters navon2023equivariant; kofinas2024graph. A parallel research direction focuses on building neural networks that process weights as inputs. Neural Functional Transformers zhou2023nft and permutation-equivariant neural functionals zhou2023permutation construct architectures that can extract information from network parameters while respecting their symmetries. Methods operating on Low-Rank Adaptation (LoRA) weights lim2024learning develop GL-equivariant networks to process low-rank weight spaces of fine-tuned models.
Our work differs fundamentally from these weight-processing approaches: Rather than building encoders or networks that consume weights as inputs, we investigate whether optimization-derived weights can directly serve as representations without additional encoding steps. This distinction positions our contribution as exploring the intrinsic representational quality of weight spaces rather than developing architectures for weight-space processing.
@@
Generative modeling in weight space has seen significant progress, with diffusion models now capable of synthesizing functional neural networks. Weight-space generation with diffusion models wang2024neural; wang2025scaling; peebles2023learning has been explored for generating models for image classification. With a different focus, our work explore neural neural field weights as representation of data that support high-quality generation and encode semantic structure.
2.2 Implicit Neural Representations
Implicit Neural Representations (INRs), also known as neural fields, are continuous functions parameterized by neural networks that map coordinates to signal values. INRs represent signals as continuous functions $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, where a neural network maps $n$-dimensional coordinates to $m$-dimensional quantities. This paradigm enables resolution-independent, memory-efficient representations of complex signals.
The versatility of INRs has been extensively documented xie2022neural; essakine2024where. These representations are fundamentally data-agnostic, employing identical network architectures across diverse modalities including 1D audio, 2D images, 3D shapes, and even 4D spatiotemporal data.
Beyond single-instance fitting, generalizable INRs have been proposed to learn priors across datasets through approaches based on autoencoders park2019deepsdf, generative adversarial networks (GANs) anokhin2021cips; karras2021alias; chan2020pigan and shared layers vyas2024learning. The GAN-based works could be seen as extensions of the StyleGAN karras2019style paradigm into the realm of neural fields. They generate different instances by modulating an MLP trunk, upon which we build as base network.
Because INRs parameterize data as neural network functions, the weights offer a direct pathway to data representation. This perspective has practical applications in compression, with methods dupont2021coin; gordon2024doh demonstrating competitive compression ratios by storing quantized network parameters instead of raw data. However, whether the collection of weights could encode semantic structure remains an open question. Another line of work employs hypernetworks klocek2019hypernetwork and transformers chen2022transformers to predict INR weights from input data via learned mappings as way of data generation. In contrast, we investigate whether gradient-descent-optimized weights can directly serve as meaningful representations. HyperDiffusion erkocc2023hyperdiffusion train a diffusion transformer to generate neural field weights as mean of synthesizing 3D shapes and 4D animated shapes. Our work build on this work to further explore what ensures successful weight space generative modeling as well as explore semantic structures in neural field weights.
@@ 3.1 Weight Space Representation
where $\mathbf{p}$ is a spatial coordinate. The architecture consists of a Fourier Feature layer followed by two linear layers. We evaluate this approach as a baseline, to encourage consistency, all MLPs share the same random initialization across different instances as is done in erkocc2023hyperdiffusion.
Encouraged by the observation that LoRA weights converge to a certain a subspace hu2022lora, which indicates that the base network enforce a certain structure on the space of LoRA weights, we fine tune a pre-trained base model using Low-Rank Adaptation (LoRA). For each instance $\mathbf{x}_{i}$, we optimize LoRA parameters $\boldsymbol{\phi}_{i}=\{\mathbf{A}_{i}^{l},\mathbf{B}_{i}^{l}\}_{l=1}^{L}$ across $L$ layers while keeping the base weights frozen. That is, we solve
@@ 3.2 Multiplicative LoRA
We employ multiplicative LoRA rather than the standard additive formulation. Standard LoRA hu2022lora adapts a pre-trained weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ through additive low-rank updates $\mathbf{W}^{\prime}=\mathbf{W}+\mathbf{B}\mathbf{A}$ where $\mathbf{A}\in\mathbb{R}^{r\times d_{\text{in}}}$ and $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ are low-rank matrices with rank $r\ll\min(d_{\text{in}},d_{\text{out}})$. We introduce a multiplicative formulation that applies weight updates through elementwise multiplication as
@@
where $\odot$ denotes elementwise multiplication. This formulation enables more effective modulation of features, analogous to successful modulation techniques in generative neural fields anokhin2021cips; karras2021alias; chan2020pigan. In our experiments, we demonstrate that this design is a critical to obtaining good weight space structure and performance on reconstruction, generation and discriminative tasks.
@@ 3.3 Base Model Architecture and Training
For LoRA-based representations, we require a strong base model that captures transferable features across the data distribution. We adopt a coordinate-based neural field architecture with multiplicative weight modulation, a design found across multiple generative neural field works anokhin2021cips; karras2021alias; chan2020pigan. The network consists of an MLP-based trunk, where variations across instances are injected through multiplicative weight modulation. This modulation mechanism naturally aligns with our multiplicative LoRA formulation, and the architecture is applicable, but not limited, to 2D and 3D data.
We train the base model using the variational autodecoder paradigm park2019deepsdf. This training scheme is desirable because it requires no encoder design, aligning with the data-agnostic quality of INRs. Given a dataset $\{\mathbf{x}_{i}\}_{i=1}^{N}$, we jointly optimize the network parameters $\boldsymbol{\theta}$ and per-instance latent codes $\{\mathbf{z}_{i}\}_{i=1}^{N}$ by solving
@@
3.4 Asymmetric Mask for Permutation Symmetry
Permutation symmetry refers to the invariance of network functions under neuron reordering, which causes functionally identical networks to occupy vastly different locations in weight space zhao2025symmetry; gao2024revisiting; navon2023equivariant. Although LoRA weights converge to a certain subspace hu2022lora, permutation symmetry remains present even within LoRA parameterizations, affecting both additive and multiplicative formulations. In additive LoRA, the operation $\mathbf{B}\mathbf{A}$ can be viewed as a two layer network: $\mathbf{A}$ acts as an encoder compressing to $r$ hidden activations, and $\mathbf{B}$ acts as a decoder expanding back. Permuting these $r$ hidden dimensions preserves the function, inducing an $r!$-fold symmetry. In multiplicative LoRA, the adapted weight matrix decomposes into a sum of terms, each being the base weight matrix scaled by diagonal matrices. Permuting the rank dimensions in this sum preserves the output due to commutative addition, again leading to permutation symmetry. We provide formal proofs for both cases in the appendix. Notably, multiplicative LoRA weights are aligned with base network channels once permutation symmetry is eliminated, see Corollary 2.2.
To address this symmetry, we investigate the asymmetric masking technique of lim2024empirical, designed to eliminate permutation symmetry during optimization. For each layer, we randomly freeze $\sqrt{d_{\text{out}}}$ entries in each row of the weight matrix, where $d_{\text{out}}$ is the out dimension. The frozen entries and their positions are shared across different runs. The values of the frozen entries are determined as follows. Suppose entry $\mathbf{W}_{ij}$ is chosen to be fixed. For the standalone MLP and additive LoRA, we initialize the frozen entries with higher variance: $\mathbf{W}_{ij}\sim\mathcal{N}(0,\kappa\mathbf{I})$. with the other weights being initialized with $\mathcal{N}(0,\mathbf{I})$. In the case of multiplicative LoRA, we simply zero out the frozen entries: $\mathbf{W}_{ij}\leftarrow 0$. This is natural for the multiplicative formulation.
While asymmetric masking was originally developed for standalone MLPs and can be applied to all three parameterizations, it proves most effective for multiplicative LoRA in the neural field domain. For standalone MLPs and additive LoRA, the technique requires large variance $\kappa$ for the frozen entries to break symmetry effectively lim2024empirical. However, this is problematic for neural fields, which synthesize signals through additive composition. In this regime, fixing certain weights to large magnitudes creates entanglement: other weights must compensate by canceling out these fixed signals, leading to difficult optimization landscapes. Multiplicative LoRA avoids this issue by zeroing out frozen entries, effectively removing certain signal components rather than forcing compensation. This approach aligns naturally with the multiplicative structure and avoids weight entanglement. We validate this advantage empirically in our experiments.

### v2

# Weight Space Representation Learning via Neural Field Adaptation

@@ Abstract
In this work, we investigate the potential of weights to serve as effective representations, focusing on neural fields. Our key insight is that constraining the optimization space through a pre-trained base model and multiplicative low-rank adaptation (mLoRA) can induce structure in weight space. Across reconstruction, generation, and analysis tasks on 2D and 3D data, we find that mLoRA weights achieve high representation quality while exhibiting distinctiveness and semantic structure. When used with latent diffusion models, mLoRA weights enable higher-quality generation than existing weight-space methods.
@@ 1 Introduction
Neural network weights have traditionally been viewed as opaque byproducts of optimization, high-dimensional vectors that encode learned functions but resist interpretation or manipulation. This perspective has begun to shift with recent advances in weight space learning, where researchers have demonstrated that network parameters can be merged [42, 35, 25], generated [9, 30], or used as inputs to other networks [46, 17, 24]. A fundamental question nonetheless remains largely unexplored: Can neural network weights themselves serve as meaningful representations for data?
@@
We investigate this question in the context of implicit neural representations (INRs), where neural networks are trained to overfit individual samples by mapping coordinates to values. INRs have proven to be versatile, capable of encoding diverse data modalities within a unified architecture [41]. Since INRs inherently encode signals as network parameters, using these weights as representations is a natural next step. However, neural network weights are known to be ambiguous by nature, for example because neuron permutations and scaling can leave the network function unchanged [44]; different random initializations may yield vastly different parameter configurations yet functionally identical models. In other words, functionally identical networks can be arbitrarily far in weight space [44], making the distribution multi-modal and difficult to learn, challenging the use of network weights as data representations.
@@
Figure 1: LoRA based weight space representation with neural fields. Given an input coordinate $\mathbf{p}\in\mathbb{R}^{n}$, a base neural field is adapted via mLoRA weights $\boldsymbol{\phi}_{i}$ to produce signal values $\mathbf{v}_{\mathbf{p}}\in\mathbb{R}^{m}$, each weight representing an instance. The mLoRA weights themselves form structured representations in weight space, enabling diverse applications.
@@
Our key insight is that constraining the network weights of different samples by introducing appropriate inductive biases allows us to transform these chaotic parameters into organized, semantic representations. To this end, as illustrated in Figure 1, we propose to use Low-Rank Adaptation (LoRA) [18] within a pre-trained base neural field to create structured weight space representations. This design is motivated by two key properties of LoRA: First, adaptation through LoRA constrains the weight updates to lie within a low-dimensional subspace defined by the base model. Hu et al. [18] demonstrate this through subspace similarity analysis, showing that adaptations at different ranks share common singular vector directions, suggesting the existence of a meaningful low-rank adaptation subspace. Second, the low-rank constraint inherently reduces the dimensionality of the weight space representation, mitigating the curse of dimensionality that would otherwise hinder learning in high-dimensional parameter spaces.
@@
We find that the standard additive LoRA formulation is insufficient for weight space learning in the context of neural fields. Instead, we introduce multiplicative LoRA (mLoRA), where weight updates are applied through element-wise multiplication rather than addition. This multiplicative formulation naturally aligns with modulation mechanisms in generative neural fields [1, 20, 3], where features are composed through multiplicative interactions, enabling effective weight space learning.
@@
We validate our approach across multiple data modalities and tasks. First, we establish that LoRA-based weight representations achieve lower reconstruction error, greater consistency under different initializations, and better linear mode connectivity than standalone MLP weights. Second, we demonstrate that these structured weight spaces support generative modeling, with diffusion models trained on multiplicative LoRA weights outperforming prior attempts in weight space neural field generation. Finally, through evaluation on discriminative tasks (i.e. classification and clustering), we confirm that the weight space structure correlates with semantic properties of the encoded data.
@@
1. We demonstrate that independently optimized neural network weights, when properly constrained, can serve as effective data representations that capture semantic structure.
2. We introduce multiplicative LoRA for neural fields, which we show provides superior representation quality compared to standard additive LoRA and standalone weight parameterizations.
3. We validate weight space representations across diverse tasks: Reconstruction, generation, and classification, establishing their viability as a representation paradigm.
@@ 2.1 Weight Space Learning
The treatment of neural network weights as learnable representations has emerged as a distinct research direction, progressing from early hypernetwork approaches to sophisticated weight-space generative models. Early work demonstrated that network parameters could be generated by auxiliary networks [15], though these methods suffered from prohibitive memory overhead when scaling to modern architectures [39]. The fundamental challenge of weight space manipulation stems from permutation symmetry: Functionally identical networks can have vastly different parameter configurations due to neuron reordering [13, 27].
@@
Recent advances have addressed these challenges through multiple strategies. Model merging techniques leverage optimal transport and activation matching to align neurons before parameter fusion [34, 28], while equivariant architectures explicitly respect weight-space symmetries when processing network parameters [27, 23]. A parallel research direction focuses on building neural networks that process weights as inputs. Neural Functional Transformers [46] and permutation-equivariant neural functionals [45] construct architectures that can extract information from network parameters while respecting their symmetries. Methods operating on Low-Rank Adaptation (LoRA) weights [24] develop GL-equivariant networks to process low-rank weight spaces of fine-tuned models. Our work takes a different approach: rather than building model-agnostic weight encoders that process weights as external data [46, 45, 27, 24], we focus on enforcing structure directly on the weight space itself, through the choice of adaptation mechanism (multiplicative LoRA) and symmetry-breaking constraints (asymmetric masking). This makes the weights serve as effective representations without requiring additional encoding steps.
@@
Generative modeling in weight space has seen significant progress, with diffusion models now capable of synthesizing functional neural networks. Weight-space generation with diffusion models [40, 39, 30] has been explored for generating models for image classification. Dravid et al. [6] show that LoRA weights of diffusion models fine-tuned on human identities form an interpretable linear subspace, enabling semantic editing and sampling via PCA. With a different focus, our work explores neural field weights as representations of data that support high-quality generation and encode semantic structure.
2.2 Implicit Neural Representation
. Implicit Neural Representations (INRs), or neural fields, parameterize signals as continuous functions $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ via neural networks, enabling resolution-independent, modality-agnostic representations across images, shapes, and spatiotemporal data [41, 10]. Generalizable INRs learn dataset-level priors through autodecoders [29], GANs with modulated MLP trunks [1, 20, 3, 21], and shared-layer schemes [38]. Since INR weights directly parameterize data, they have been explored for compression [7, 14], hypernetwork-based generation [22, 5], and meta-learned modulation representations [8]. In contrast to these approaches, we investigate whether independently optimized weights can directly serve as meaningful representations. Closely related to our work, HyperDiffusion [9] trains a diffusion transformer over neural field weights to synthesize 3D and 4D shapes. We build on this line of work, investigating how the choice of adaptation mechanism and symmetry-breaking constraints affect weight space generation quality and semantic structure. A more detailed discussion is provided in the supplementary material (Section 7).
@@ 3.1 Weight Space Representation
where $\mathbf{p}$ is a spatial coordinate. The architecture consists of a Fourier Feature layer followed by two linear layers. We evaluate this approach as a baseline, to encourage consistency, all MLPs share the same random initialization across different instances as is done in [9].
Encouraged by the observation that LoRA weights converge to a certain a subspace [18], which indicates that the base network enforce a certain structure on the space of LoRA weights, we fine tune a pre-trained base model using Low-Rank Adaptation (LoRA). For each instance $\mathbf{x}_{i}$, we optimize LoRA parameters $\boldsymbol{\phi}_{i}=\{\mathbf{A}_{i}^{l},\mathbf{B}_{i}^{l}\}_{l=1}^{L}$ across $L$ layers while keeping the base weights frozen. That is, we solve
@@ 3.2 Multiplicative LoRA
We employ multiplicative LoRA rather than the standard additive formulation. Standard LoRA [18] adapts a pre-trained weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ through additive low-rank updates $\mathbf{W}^{\prime}=\mathbf{W}+\mathbf{B}\mathbf{A}$ where $\mathbf{A}\in\mathbb{R}^{r\times d_{\text{in}}}$ and $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ are low-rank matrices with rank $r\ll\min(d_{\text{in}},d_{\text{out}})$. We introduce a multiplicative formulation that applies weight updates through elementwise multiplication as
@@
where $\odot$ denotes elementwise multiplication. This formulation enables more effective modulation of features, analogous to successful modulation techniques in generative neural fields [1, 20, 3]. In our experiments, we demonstrate that this design is critical to obtaining good weight space structure and performance on reconstruction, generation, and discriminative tasks.
We hypothesize that the advantage of multiplicative over additive LoRA is related to feature entanglement in neural fields. INRs synthesize signals through additive composition: linear layers combine basis functions while activation functions generate harmonics [43]. This additive synthesis inherently creates entangled representations. Additive LoRA exacerbates this by injecting new signal components into the already-entangled mixture, making the weight space harder to structure. In contrast, multiplicative LoRA scales existing features rather than injecting new ones, preserving channel structure and avoiding further entanglement. This aligns with Corollary 2.2, which shows that once permutation symmetry is eliminated, mLoRA weights are fully aligned with the base network’s channel axes.
@@ 3.3 Base Model Architecture and Training
For LoRA-based representations, we require a strong base model that captures transferable features across the data distribution. We adopt a coordinate-based neural field architecture with multiplicative weight modulation, a design found across multiple generative neural field works [1, 20, 3]. The network consists of an MLP-based trunk, where variations across instances are injected through multiplicative weight modulation. This modulation mechanism naturally aligns with our multiplicative LoRA formulation, and the architecture is applicable, but not limited, to 2D and 3D data.
We train the base model using the variational autodecoder paradigm [29]. This training scheme is desirable because it requires no encoder design, aligning with the data-agnostic quality of INRs. Given a dataset $\{\mathbf{x}_{i}\}_{i=1}^{N}$, we jointly optimize the network parameters $\boldsymbol{\theta}$ and per-instance latent codes $\{\mathbf{z}_{i}\}_{i=1}^{N}$ by solving
@@
3.4 Addressing Permutation Symmetry
Permutation symmetry refers to the invariance of network functions under neuron reordering, which causes functionally identical networks to occupy vastly different locations in weight space [44, 13, 27]. Permutation symmetry creates ambiguity, making the distribution multi-modal and difficult to learn. Removing this symmetry collapses these modes into a canonical representation, yielding a smoother, more structured weight space that could be effectively modeled.
Permutation symmetry has two distinct sources in our setting. External symmetries arise from permutations of the base network neurons; these are fully eliminated by fixing a single shared base model across all instances. Internal symmetries arise within the LoRA factors themselves: Both additive and multiplicative formulations permit permuting the $r$ rank dimensions without changing the represented function (we provide formal proofs in the supplementary material). Moreover, any invertible matrix $\mathbf{G}\in GL(r)$ gives $(\mathbf{A}\mathbf{G})(\mathbf{G}^{-1}\mathbf{B})=\mathbf{A}\mathbf{B}$ [24], meaning the weight space has a $GL(r)$-fold equivalence class for each represented function.
To address internal symmetry, we investigate the asymmetric masking technique of [25], applied to all LoRA $\mathbf{A}$ matrices across all layers. For each layer, we randomly freeze $\sqrt{d_{\text{out}}}$ entries in each row of $\mathbf{A}$, where $d_{\text{out}}$ is the output dimension. The frozen positions are shared across all instances and training runs. For the standalone MLP and additive LoRA, frozen entries are initialized with higher variance: $\mathbf{W}_{ij}\sim\mathcal{N}(0,\kappa\mathbf{I})$, with other weights initialized with $\mathcal{N}(0,\mathbf{I})$. For multiplicative LoRA, we zero out the frozen entries: $\mathbf{A}_{ij}\leftarrow 0$, which is natural since it removes the corresponding rank component’s contribution.
While asymmetric masking can be applied to all three parameterizations, it proves most effective for multiplicative LoRA in the neural field domain. For standalone MLPs and additive LoRA, the technique requires large variance $\kappa$ for the frozen entries to break symmetry effectively [25]. However, this is problematic for neural fields, which synthesize signals through additive composition: fixing certain weights to large magnitudes creates entanglement where other weights must compensate by canceling these fixed signals, leading to difficult optimization landscapes. Multiplicative LoRA avoids this issue by zeroing out frozen entries, effectively gating off certain rank components rather than forcing compensation. This approach aligns naturally with the multiplicative structure and avoids weight entanglement. We validate this advantage empirically in our experiments.